OpenStudy (trojanpoem):

Maxima , minima :

OpenStudy (trojanpoem):

\[f(x) = \frac{ x^2 + 9 }{ x } \]

OpenStudy (trojanpoem):

Find the upwards and downwards concavity and the turning points if they exists.

OpenStudy (trojanpoem):

@Michele_Laino

OpenStudy (michele_laino):

we have to compute the first derivative of your function f(x)

OpenStudy (trojanpoem):

I did , the whole problem is that the turining point is (0 , undefined ) which is senseless for me.

OpenStudy (michele_laino):

I got this: \[\Large f'\left( x \right) = \frac{{2x \cdot x - \left( {{x^2} + 9} \right) \cdot 1}}{{{x^2}}} = \frac{{{x^2} - 9}}{{{x^2}}}\]

OpenStudy (michele_laino):

now we have to solve this inequality: \[\frac{{{x^2} - 9}}{{{x^2}}} \geqslant 0\]

OpenStudy (michele_laino):

since at those points at which f' is positive, the function f is increasing

OpenStudy (michele_laino):

what is the solution of the inequality above?

OpenStudy (trojanpoem):

Concavity not maxima ,

OpenStudy (michele_laino):

yes! I know it

OpenStudy (trojanpoem):

x^2 - 9 > x^2 -9 > 0

OpenStudy (michele_laino):

the first derivative is positive, when the numerator only is positive, since the denominator is always positive in all points x such that x is different from zero

OpenStudy (michele_laino):

so we have to solve this equivalent inequality: \[{x^2} - 9 \geqslant 0\]

OpenStudy (trojanpoem):

I got laggy and disconnected.

OpenStudy (trojanpoem):

3 > x > -3

OpenStudy (michele_laino):

the solution is: \[x \geqslant 3,\quad x \leqslant - 3\] so, we can plot your solution as below: